OpenStudy (anonymous):

solve the system of linear equation with the elimination method: 2x+y=1 3x+6y+30

OpenStudy (anonymous):

2x+y=1 <=> -12x-6y=-6 add that to the second equation 3x-12x-6y+6y+30=0 <=>(eliminate y) => -9x+30=0 => x=10/3 Plug that x in the first equation, 20/3+y=1 => y=-17/3

OpenStudy (anonymous):

2x+y=1 ..........eq(1) 3x+6y+30.............eq(2) multiply eq(1) by 3 and eq(2) by 2 then u wll get 6x+3y=3 ..........eq(3) 6x+12y=60 ..........eq(4) then subtract eq(3) from eq(4)..then 6x will cancel 6x then left with 9y=57 y =57/9 then take y and replace @ eq(1) u will get.... 2x=1-(57/3) then divide both sides by 2.... u will get x=(1-(57/3))/2